OpenStudy (anonymous):

Just factored and got this (2x+2x) what do i do with this?

OpenStudy (anonymous):

@amistre64 @ash2326 @Hero

OpenStudy (anonymous):

hey

OpenStudy (anonymous):

@Hero can you help?

OpenStudy (anonymous):

add, since they have the same variable

hero (hero):

Just add them together bro

OpenStudy (anonymous):

lol

hero (hero):

Or you could do something weird like this: (2x + 2x) = 2(x + x) = 2(1 + 1)x = 2(2)x = 4x

OpenStudy (anonymous):

\[4x ^{2}+6x+2\] I got it from this

hero (hero):

Suggestion bro, always post the original problem you're working on

OpenStudy (anonymous):

ok mam

OpenStudy (anonymous):

\[(4x+2)(x+1) \]

OpenStudy (anonymous):

Well im a sis not a bro

hero (hero):

Lynncake that's not factored all the way

OpenStudy (anonymous):

you get two values of x \[4x+2 =0\rightarrow 4x=-2 \rightarrow x=\frac{ -1 }{ 2 }\] \[x+1=0 \rightarrow x=-1\]

OpenStudy (anonymous):

i got \[(2x+2) and (2x+2x)\]

hero (hero):

See, that's what I mean @tomtom777, always post the original question you're working on.

OpenStudy (anonymous):

that's not equal to \[4x^2+6x+2\]

hero (hero):

That's not even close to right

OpenStudy (anonymous):

yours equates to\[8x^2 + 8x \]

OpenStudy (anonymous):

see mine @tomtom777

hero (hero):

@Lynncake, you did not factor completely.

hero (hero):

I can post the full factorization if you don't believe me.

OpenStudy (anonymous):

im factoring my real problem is lol this\[\frac{ 4x ^{2}+6x+2 }{ 4x ^{2}-5x-6 }\]

OpenStudy (anonymous):

i need to find the undefined

hero (hero):

See, that's what you should have posted from the beginning.

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

lol i know

OpenStudy (anonymous):

\[\frac{ (4x+2) (x+1) }{ (4x+3)(x-2) }\]is this right? @Hero

hero (hero):

For the numerator, the factorization is: \[4x^2 + 6x + 2 \\ = 2(2x^2 + 3x + 1) \\ = 2(2x^2 + 2x + x + 1) \\ = 2(2x(x + 1) + 1(x + 1)) \\ = 2((x + 1)(2x + 1)) \\= 2(2x + 1)(x + 1)\]

OpenStudy (anonymous):